OpenStudy (anonymous):

using the point-slope form how do i write an equation for a line containing the points (1,-3) and (-5,2)

OpenStudy (anonymous):

find the slope. (2-(-3))/(-5-1)=5/-6=-5/6=m Plug this into y=mx+b, along with either of the ordered pairs for y and x. This will give you b. The final equation will be y=-5/6m+b (I'm too lazy to solve for b)